# SOC-DGL: Social Interaction Behavior Inspired Dual Graph Learning Framework for Drug-Target Interaction Identification

### v1

@@ 3.2.6 Imbalanced Processing
To address the issue of class imbalance between positive and negative samples in the model, this study designs a Weighted Binary Cross-Entropy Loss Function with Reduced Regative Samples (RLF), as shown in the following formula:
$\mathcal{L}=-\frac{1}{n_{d}\cdot n_{t}}(\varpi\frac{\mid y^{-}\mid}{\mid y^{+}\mid}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (14)
where $n_{d}$ is the number of drugs, $n_{t}$ is the number of targets, $|y^{+}|$ represents the number of positive samples, and $|y^{-}|$ represents the number of negative samples. $h_{i}^{*}$ represents the model’s predicted probability of interaction between drug $i$ and target $j$. This novel loss function is built upon the existing weighted cross-entropy loss function, with the introduction of an adjustable parameter $\varpi$, which effectively reduces the weights of negative samples while keeping the weights of positive samples unchanged.
Table 1: SOC-DGL performance in the “warm-start” scenario on four benchmark datasets.
Dataset KIBA dataset Davis dataset BindingDB dataset DrugBank dataset
Situation balanced imbalanced balanced imbalanced balanced imbalanced balanced imbalanced
AUROC 0.9761 $\pm$ 0.0023 0.9364 $\pm$ 0.0036 0.9467 $\pm$ 0.0074 0.9192 $\pm$ 0.0045 0.9894 $\pm$ 0.0032 0.9790 $\pm$ 0.0034 0.9735 $\pm$ 0.0023 0.9295 $\pm$ 0.0076
AUPR 0.9805 $\pm$ 0.0018 0.7473 $\pm$ 0.0123 0.9673 $\pm$ 0.0048 0.7655 $\pm$ 0.0134 0.9888 $\pm$ 0.0036 0.8705 $\pm$ 0.0146 0.9774 $\pm$ 0.0019 0.7654 $\pm$ 0.0126
F1_score 0.9350 $\pm$ 0.0041 0.6908 $\pm$ 0.0068 0.9038 $\pm$ 0.0103 0.6918 $\pm$ 0.0129 0.9641 $\pm$ 0.0075 0.8057 $\pm$ 0.0177 0.9286 $\pm$ 0.0027 0.7211 $\pm$ 0.0104
ACC 0.9268 $\pm$ 0.0049 0.9207 $\pm$ 0.0026 0.8816 $\pm$ 0.0131 0.9134 $\pm$ 0.0030 0.9627 $\pm$ 0.0077 0.9594 $\pm$ 0.0041 0.9267 $\pm$ 0.0026 0.9508 $\pm$ 0.0015
Recall 0.9373 $\pm$ 0.0093 0.7103 $\pm$ 0.0182 0.9097 $\pm$ 0.0200 0.6642 $\pm$ 0.0338 0.9677 $\pm$ 0.0098 0.8406 $\pm$ 0.0246 0.9208 $\pm$ 0.0136 0.6567 $\pm$ 0.0210
Specificity 0.9133 $\pm$ 0.0148 0.9506 $\pm$ 0.0049 0.8372 $\pm$ 0.0362 0.9561 $\pm$ 0.0066 0.9574 $\pm$ 0.0107 0.9726 $\pm$ 0.0047 0.9331 $\pm$ 0.0130 0.9824 $\pm$ 0.0022
Precision 0.9328 $\pm$ 0.0100 0.6730 $\pm$ 0.0177 0.8985 $\pm$ 0.0189 0.7238 $\pm$ 0.0230 0.9606 $\pm$ 0.0097 0.7744 $\pm$ 0.0304 0.9369 $\pm$ 0.0109 0.8004 $\pm$ 0.0174
Figure 2: Performance comparison of SOC-DGL with baseline methods on the KIBA dataset. (A) ROC and PR curves on balanced datasets and (B) imbalanced datasets. (C) F1 score, ACC, recall, specificity, and precision on balanced datasets and (D) imbalanced datasets.
@@ 4.1 Experimental design overview
In the ”warm-start” scenario, this study designed validation experiments for both balanced datasets (positive:negative samples = 1:1) and imbalanced datasets (positive:negative samples = 1:10). In most real-world tasks, the ratio of positive to negative samples is typically imbalanced. In drug-target association prediction, known associations (positive samples) were generally far fewer than unknown associations (negative samples). Evaluating with imbalanced datasets better reflected SOC-DGL’s performance in real-world settings. Conversely, balancing the positive and negative samples helped mitigate model bias towards predicting negative samples, providing a more balanced assessment of model performance.
To minimize the impact of data variability on the results, this study employed 10-fold cross-validation to assess the model’s predictive performance. First, the entire dataset (positive and negative samples) was randomly divided into 10 subsets of approximately equal size. This process was repeated 10 times, with each time selecting a different subset as the test set and the other 9 subsets as the training set. Initially, the test set data was masked, and then the model was trained on the remaining 9 subsets. The masked test set values were predicted for DTI, and the predicted values were compared with the actual DTI values (before masking) using the AUROC, AUPR, F1_score, Accuracy (ACC), Recall, Specificity, and Precision metrics.
In this study, in the “warm-start” scenario, comprehensive experimental validation was conducted on four representative benchmark datasets to assess the performance of SOC-DGL. For other specific experiments, the study primarily chose to conduct validation on the KIBA dataset due to its moderate data size and comprehensive information, which provided accurate support for model training and evaluation.
This study conducted experiments using PyTorch 1.11.0 on an Nvidia GeForce RTX 3060 GPU, and employed the Xavier initialization strategy to prevent gradient explosion and vanishing during the training process.
4.2 “Warm-start” experiments
In this study, “Warm-start” experiments were conducted on four benchmark datasets under both balanced and imbalanced settings. Multiple runs were performed on each dataset with optimized parameter configurations, and average metrics with standard deviations were reported, as shown in Tab. 1. The results demonstrated that SOC-DGL achieved high predictive accuracy in balanced settings, effectively distinguishing known DTIs, and performed well in the “warm-start” scenario, validating application in DTI prediction and drug discovery.
@@
To assess SOC-DGL’s effectiveness, comparative experiments were conducted with four leading DTI prediction methods: GraphDTA [27], LRSpNM [28], MLMC [29], MULGA [30], and MMDG-DTI [31], representing the state-of-the-art. Detailed comparisons were available in Appendix Section C. SOC-DGL showed significant improvements on both balanced and imbalanced datasets, demonstrating its broad applicability and outperformance over current methods.
As shown in Fig. 2(A) and Fig. 2(B), SOC-DGL consistently outperformed all baseline methods on both balanced and imbalanced datasets. It was clearly evident that MMDG-DTI consistently performed just slightly behind SOC-DGL on the KIBA balanced dataset. MLMC also showed strong performance, leveraging multi-view learning and matrix completion techniques. In contrast, LRSpNM, based on low-rank matrix approximation, performed the worst across all datasets, likely due to DPI matrix sparsity and inaccurate similarity measures. On the KIBA balanced dataset, SOC-DGL’s AUROC and AUPR were 1.8% and 1.6% higher compared to the current best-performing model, respectively, and 0.4% and 0.1% higher in the imbalanced setting. MLMC and MULGA performed slightly worse than SOC-DGL, with MULGA underperforming in imbalanced settings due to inadequate attention to homogeneous association information. However, in other datasets, MULGA’s multi-view learning approach combined with graph autoencoders ranked second to SOC-DGL, suggesting lower generalizability. Appendix Tab. I further confirmed SOC-DGL’s superiority across datasets. The F1-score, ACC, Recall, Specificity, and Precision results in Fig. 2(C) and Fig. 2(D) further validated SOC-DGL’s consistent outperformance.
Figure 3: Performance scores of SOC-DGL in ”cold-start” scenarios for drugs and targets, evaluated on various metrics. Black dots represent randomly sampled data points corresponding to indicator values of these metrics.
@@
In the “cold-start” experiments on the KIBA dataset, two scenarios were designed: “cold-start for drugs” and “cold-start for targets,” where the drug-target affinity matrix was systematically sampled. In the “cold-start for targets” scenario, drugs linked to specific targets were removed, treating each target as new, with the removed data serving as the test set. In the “cold-start for drugs” scenario, drugs linked to specific targets were removed, and the corresponding test set consisted of the removed drugs and their targets.
@@
The KIBA dataset, comprising 1,720 drugs and 220 targets, provided the foundation for “Cold-start” experiments. A total of 56 independent cold-start experiments were conducted, involving the random selection of 34 drugs and 22 targets to simulate real-world scenarios where new drugs or targets lacked historical interaction data. This was particularly relevant in drug discovery, where novel compounds or under-explored proteins often lacked prior interaction data.
@@
Fig. 3 presented the results. In the “cold-start for drugs” scenario, SOC-DGL showed stable performance across different drugs, with an average AUROC of 0.9758 and AUPR of 0.9803, indicating strong generalization ability for new drugs. Similarly, in the “cold-start for targets” scenario, SOC-DGL demonstrated consistent performance across different targets. These results confirmed its robustness in predicting drug-target interactions for unseen drugs or targets, making it a valuable tool for practical drug development and target discovery.
4.4 Analysis of the Even-Polynomial Graph Filter in EDGL
To analyze the importance of the even-polynomial graph filter in EDGL for indirectly extracting higher-order homogeneous information between drugs and targets, this study compared it with the odd-polynomial graph filter for indirectly extracting higher-order heterogeneous information and with the self-attention mechanism under the “warm start” scenario, while keeping all other modules unchanged. Specifically, this study evaluated the effectiveness of higher-order similarity information enhancement by the even-polynomial graph filter (Homogeneous_prop), higher-order interaction information enhancement by the odd-polynomial graph filter (Heterogeneous_prop), and self-information enhancement by the self-attention mechanism (Self_attention_prop) using various metrics to comprehensively measure their performance.
As shown in Fig. 4(A), Homogeneous_prop outperformed Heterogeneous_prop and Self_attention_prop, particularly in AUROC and AUPR, excelling in DTI prediction and handling imbalanced data by capturing complex indirect relationships. While Heterogeneous_prop was effective for direct connections, it struggled with indirect ones. Self_attention_prop, despite its node-weighting capability, performed weaker overall in managing complex interactions. Homogeneous_prop showed a trade-off with slightly lower Recall, indicating reduced sensitivity to some positives. These results highlighted the superior ability of Homogeneous_prop to capture indirect interactions and improve prediction accuracy.
4.5 Ablation Study on Diversity Graph Learning
To evaluate the impact of different graph learning modules, three SOC-DGL variants were tested: (1) ADGL-Only, focusing on global associations, (2) EDGL-Only, emphasizing higher-order similarities, and (3) Fusion-Without, combining the two modules without feature fusion residual connection to assess the role of feature fusion.
As shown in Fig. 4 (B), SOC-DGL outperformed the variants on most metrics, particularly in AUROC and AUPR. On balanced datasets, SOC-DGL achieved an AUROC of 0.9761, surpassing ADGL-Only (0.9660) and EDGL-Only (0.9703), demonstrating the clear benefit of combining ADGL and EDGL for capturing complex DTIs. Similarly, SOC-DGL’s AUPR of 0.9805 exceeded ADGL-Only (0.9707) and EDGL-Only (0.9690). While Fusion-Without showed comparable performance in some cases, it fell short overall, indicating that the residual fusion method, which effectively balanced the importance of different DTI patterns, more effectively identified and integrated complex interaction pathways, thereby improving model generalization.
Figure 4: Experimental validation conducted on the KIBA dataset. (A) Experimental evaluation of Homogeneous_prop, Heterogeneous_prop, and Self_attention_prop enhancement components in SOC-DGL for DTI prediction. (B) Ablation study of different components within the SOC-DGL model.
@@ 4.5 Ablation Study on Diversity Graph Learning
4.6 Validation Experiment on Imbalanced Loss Function
In DTI prediction, there was often an imbalance between positive and negative samples, which could affect model performance, causing it to favor the majority class while neglecting the minority. This experiment evaluated the impact of various loss functions on model performance, focusing on those designed to address class imbalance.
Four loss functions were assessed: Standard Binary Cross-Entropy Loss (SLF), Focal Loss with Dynamic Scaling (FLF), Weighted Binary Cross-Entropy Loss (WLF), and Weighted Binary Cross-Entropy Loss with Reduced Negative Samples (RLF). Experiments were conducted in a “warm-start” scenario with imbalanced datasets, evaluating DTI prediction performance across multiple metrics. Details of these loss functions could be found in Appendix Section E.
As shown in Tab. 2, the RLF loss function consistently outperformed the others across four benchmark imbalanced datasets, demonstrating its effectiveness for imbalanced DTI prediction. Although RLF did not always yield the highest scores on individual metrics, it showed stable and robust performance. By incorporating a weighting parameter $\varpi$ to reduce the weight of negative samples while maintaining the positive sample weight, RLF addressed class imbalance while preserving the model’s discriminative ability. These results, observed across four distinct imbalanced datasets, further validated RLF’s strong generalization capability.

### v2

@@ 3.2.6 Imbalanced Processing
To address the issue of class imbalance between positive and negative samples, this study designs a Weighted Binary Cross-Entropy Loss Function with Reduced Regative Samples (RLF):
$\mathcal{L}=-\frac{1}{n_{d}\cdot n_{t}}(\varpi\frac{|y^{-}|}{|y^{+}|}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (10)
where $|y^{+}|$ and $|y^{-}|$ represent the number of positive samples and negative samples. $h_{ij}^{*}$ represents the model’s predicted probability of interaction between drug $i$ and target $j$. This loss function is built upon the existing weighted cross-entropy loss function, with the introduction of an adjustable parameter $\varpi$, which effectively reduces the weights of negative samples while keeping the weights of positive samples unchanged.
3.2.7 Parameter Setting
Experiments were conducted on four benchmark datasets under balanced or imbalanced setting. For the balanced KIBA dataset, we set the learning rate to 0.00003, neighbor aggregation depth $K$ to 200, neighbor contribution ratio $\alpha$ to 0.1, residual connection weight $w$ to 0.8, and maximum training epochs to 1500. Under the imbalanced setting, an additional imbalance loss parameter $\varpi$, used to adjust the weight of negative samples, was set to 0.6. Parameter details for other datasets under balanced or imbalanced setting are provided on our GitHub repository.
@@ 4 Experiments and Results
To evaluate the effectiveness of SOC-DGL, we design five experiments set, including the performance comparison, “Cold-start” Experiments, ablation experiments (consists of three sub-experiments), hyperparameter tuning experiments (consists of three sub-experiments) and case study.
@@ 4.1 Experimental Design Overview
This study conducted validation experiments on both balanced (positive:negative = 1:1) and imbalanced (positive:negative = 1:10) datasets. Real-world tasks, particularly DTI prediction, typically involve imbalanced datasets with far fewer positive (known) than negative (unknown) samples, evaluating with imbalanced datasets provides a more realistic assessment of SOC-DGL’s performance. Balancing the dataset, however, mitigates model bias towards predicting negative samples, offering a more balanced evaluation of model effectiveness. To minimize data variability, 10-fold cross-validation was employed to evaluate model performance. During each fold, the test set was masked, and the performance is evaluated using metrics such as AUROC, AUPR, F1_score, ACC, Recall and Precision. For experiments, comprehensive validation was performed on four benchmark datasets to assess SOC-DGL’s performance. However, the KIBA dataset was primarily selected for other specific experiments due to its moderate size and comprehensive data, which ensured accurate model training and evaluation. This study conducted experiments using PyTorch 1.11.0 on an Nvidia GeForce RTX 3060 GPU, and employed the Xavier initialization strategy to prevent gradient explosion and vanishing during the training process.
Table 1: T-test Across Folds for SOC-DGL on Four Datasets under Balanced and Imbalanced Conditions.
Dataset $P(\text{KIBA})$ $P(\text{Davis})$ $P(\text{BindingDB})$ $P(\text{DrugBank})$ Balanced 0.9950 0.5959 0.2242 0.5802 Imbalanced 0.0133 0.9292 0.4479 0.0429
4.2 Performance Comparison
@@
To assess SOC-DGL’s effectiveness, comparative experiments were conducted with seven leading DTI prediction methods: GraphDTA [36], LRSpNM [37], MLMC [38], MULGA [39], MSI-DTI [4], BCMMDA [40] and MMDG-DTI [12], representing the state-of-the-art (SOTA). The Details of these methods were available in Appendix Section C.
LRSpNM 2021 0.9224 0.9382 0.8740 0.8740 0.8740 0.7730 0.8371
MLMC 2022 0.9484 0.9592 0.9049 0.8935 0.9038 0.8803 0.9059
MULGA 2023 0.9479 0.9551 0.8984 0.8773 0.9404 0.7911 0.8600
MSI-DTI 2024 0.9320 0.8370 0.7503 0.9080 0.7210 - 0.7820
MMDG-DTI 2025 0.9586 0.9646 0.9083 0.8949 0.9318 0.8483 0.8858
SOC-DGL \ 0.9761 0.9805 0.9350 0.9268 0.9373 0.9133 0.9328
LRSpNM 2021 0.9132 0.6904 0.6339 0.8998 0.6900 0.9299 0.5863
MLMC 2022 0.9329 0.7462 0.6874 0.9192 0.7084 0.9494 0.6675
MULGA 2023 0.8918 0.5556 0.5628 0.8666 0.6808 0.8934 0.4796
BCMMDA 2024 0.9270 0.8040 - - - - -
MMDG-DTI 2025 0.9245 0.7193 0.6639 0.9102 0.6998 0.9391 0.6273
SOC-DGL \ 0.9364 0.7473 0.6908 0.9207 0.7103 0.9506 0.6730
LRSpNM 2021 0.8930 0.9256 0.8644 0.8255 0.9083 0.6951 0.8250
MLMC 2022 0.9072 0.9395 0.8712 0.8712 0.9098 0.7177 0.8356
MULGA 2023 0.9158 0.9451 0.8790 0.8561 0.8539 0.8596 0.9058
MSI-DTI 2024 0.9660 0.8930 0.8570 0.9325 0.8805 - 0.8348
MMDG-DTI 2025 0.9356 0.9647 0.8905 0.8627 0.9057 0.7934 0.8758
SOC-DGL \ 0.9467 0.9673 0.9038 0.8816 0.9097 0.8372 0.8985
LRSpNM 2021 0.9108 0.7067 0.6415 0.8961 0.6348 0.9409 0.6517
MLMC 2022 0.9110 0.7253 0.6704 0.9057 0.6530 0.9492 0.6888
MULGA 2023 0.8812 0.6179 0.5710 0.8698 0.5874 0.9187 0.5556
BCMMDA 2024 0.9330 0.7800 0.7040 0.8760 0.6070 - 0.8380
MMDG-DTI 2025 0.9087 0.7787 0.7369 0.9326 0.6503 0.9805 0.8409
SOC-DGL \ 0.9192 0.7655 0.6918 0.9134 0.6642 0.9561 0.7238
LRSpNM 2021 0.8514 0.5771 0.5533 0.8706 0.5464 0.9263 0.5602
MLMC 2022 0.8797 0.9285 0.8707 0.8750 0.8237 0.9286 0.9233
MULGA 2023 0.9874 0.9859 0.9678 0.9660 0.9818 0.9490 0.9542
MMDG-DTI 2025 0.9620 0.9661 0.9181 0.9172 0.9172 0.9340 0.9340
SOC-DGL \ 0.9894 0.9888 0.9641 0.9627 0.9677 0.9574 0.9606
LRSpNM 2021 0.8481 0.5727 0.5866 0.9096 0.6383 0.9399 0.5426
MLMC 2022 0.8992 0.8260 0.7990 0.9617 0.7584 0.9844 0.8443
MULGA 2023 0.9706 0.8366 0.7188 0.9444 0.6991 0.9721 0.7395
MMDG-DTI 2025 0.9149 0.7428 0.6967 0.8980 0.6504 0.9508 0.7429
SOC-DGL \ 0.9790 0.8705 0.8057 0.9594 0.8406 0.9726 0.7744
LRSpNM 2021 0.7563 0.8206 0.7016 0.6687 0.7522 0.5790 0.6573
MLMC 2022 0.8637 0.9126 0.8371 0.8415 0.7828 0.9052 0.8995
MULGA 2023 0.9715 0.9743 0.9165 0.9141 0.9112 0.9171 0.9219
MSI-DTI 2024 0.9093 0.9136 0.8339 0.8381 0.8551 - 0.8137
MMDG-DTI 2025 0.9626 0.9658 0.9180 0.9159 0.9083 0.9240 0.9281
SOC-DGL \ 0.9735 0.9774 0.9286 0.9267 0.9208 0.9331 0.9369
LRSpNM 2021 0.7752 0.5480 0.5437 0.9257 0.4577 0.9758 0.6695
MLMC 2022 0.8594 0.7330 0.7043 0.9498 0.6181 0.9853 0.8185
MULGA 2023 0.9259 0.7179 0.6667 0.9358 0.6579 0.9659 0.6756
BCMMDA 2024 0.8560 0.8680 0.7895 0.7860 0.7920 - 0.7870
MMDG-DTI 2025 0.9051 0.6978 0.6717 0.9381 0.6522 0.9688 0.6943
SOC-DGL \ 0.9295 0.7654 0.7211 0.9508 0.6567 0.9824 0.8004
To validate the model’s robustness, this study conducted ten-fold cross-validation of SOC-DGL on four balanced and two imbalanced datasets. Paired t-tests were performed on seven metrics for each fold, followed by Fisher’s Combined Probability Test to aggregate the results into a single p-value. As shown in Table 1, the combined p-values for SOC-DGL on the four balanced and two imbalanced datasets are all greater than 0.05, except for the KIBA and DrugBank imbalanced datasets. This suggests that SOC-DGL is stable on balanced datasets and certain imbalanced datasets, demonstrating robustness, likely due to its adjustable imbalance loss function.
As shown in Table 2, the results demonstrate that the SOC-DGL model achieves robust overall performance across the four benchmark datasets. Specifically, except for the Davis dataset under the highly imbalanced data scenario where SOC-DGL performs slightly worse than MCMMDA, the model attains or surpasses SOTA results on the other datasets. This study speculate that this discrepancy primarily arises from the Davis dataset’s pronounced class imbalance, where the number of positive samples is significantly lower than negatives, posing challenges for the model to effectively learn minority class features. Nevertheless, SOC-DGL’s superior performance across diverse and relatively balanced datasets highlights the effectiveness of its graph-structured learning approach in capturing complex drug-target relationships. These results validate the soundness and practical potential of SOC-DGL.
@@
In the “cold-start” experiments, two scenarios were designed: “cold-start for drugs” and “cold-start for targets,” where the drug-target affinity matrix was systematically sampled. In the “cold-start for targets” scenario, drugs linked to specific targets were removed, treating each target as new, with the removed data serving as the test set. Similar operations were applied in the ”cold-start for drugs” scenario.
@@
The KIBA dataset contains 1,720 drugs and 220 targets, providing the foundation for the ”cold-start” experiments. A total of 1940 independent cold-start experiments were conducted, including 1720 experiments for drugs and 220 experiments for targets. These experiments were carried out by using information from other known drugs and targets to predict the selected drugs and targets, without employing 10-fold cross-validation. This setup simulates real-world scenarios where new drugs or targets lack historical interaction data, which is especially relevant in drug discovery, where novel compounds or under-explored proteins often lack prior interaction data.
Figure 2: The performance of SOC-DGL in ”cold-start” scenarios for drugs or targets, where the peak represents the degree of aggregation of results.
Figure 3: Ablation Experiments on the KIBA dataset. (A) Experimental evaluation of Homo_prop, Hetero_prop, and Self_att_prop. (B) Ablation experiment on ADGL, EDGL, and feature fusion strategies (WF and Attention) for model performance.
@@
As shown in Fig. 2, in the ”cold-start for drugs” scenario. SOC-DGL demonstrated stable performance with AUROC values ranging from [0.6, 1.0], with the most frequent results around 0.9, indicating its strong generalization ability for new drugs. In the ”cold-start for targets” scenario, the AUROC values ranged from [0.6, 0.8], which may have been influenced by the availability of target data and variability in target-specific interactions. These results confirmed SOC-DGL’s robustness in predicting DTI for unseen drugs or targets, making it a valuable tool for practical drug discovery.
@@
4.4 Ablation Experiments
4.4.1 Ablation Experiment on Even-Polynomial Graph Filter in EDGL
To analyze the importance of the even-polynomial graph filter in EDGL for indirectly extracting higher-order homogeneous information between drugs and targets, this study compared it with the odd-polynomial graph filter for extracting higher-order heterogeneous information and the self-attention mechanism for enhancing self-information under the warm-start scenario, while keeping all other modules unchanged. Specifically, this study evaluated the performance of three variants: Homo_prop (even-polynomial filter for homogeneous similarity enhancement), Hetero_prop (odd-polynomial filter for heterogeneous interaction enhancement), and Self_att_prop (self-attention mechanism for self-information enhancement). As shown in Fig. 3 (A), Homo_prop consistently achieved superior performance across all evaluation metrics. These results demonstrate the effectiveness of the even-polynomial graph filter in capturing complex high-order homogeneous relationships, enabling more accurate and robust DTI prediction.
4.4.2 Ablation Experiment on ADGL, EDGL, and Feature Fusion Strategies to Model Performance
To evaluate the impact of different graph learning modules, we tested four SOC-DGL variants: (1) ADGL, focusing on global associations, (2) EDGL, emphasizing higher-order similarities, (3) WF, where features pass through ADGL and EDGL sequentially without fusion, and (4) Attention, fusing features from ADGL and EDGL via an attention mechanism after concatenation. As shown in Fig. 3 (B), the ADGL module significantly contributes to model performance due to its extensive use of a global drug-target affinity network that captures both global DTI and individual similarity information. In contrast, the EDGL module amplifies the influence of even-hop neighbors through a higher-order strategy, enhancing structural information. Regarding feature fusion, the residual connection consistently outperforms other methods in predictive performance, offering computational advantages by avoiding matrix multiplication, which reduces parameters and training time. Biologically, the residual connection preserves the integrity of global drug-target associations and higher-order structural similarities, aligning with how functional and structural signals contribute to molecular recognition without excessive entanglement. In contrast, concatenation and attention mechanisms may obscure this interpretability by over-mixing heterogeneous information.
4.4.3 Ablation Experiment on Imbalanced Loss Function
In DTI prediction, the imbalance between positive and negative samples can affect model performance, causing it to favor the majority class. This experiment evaluated the impact of various loss functions on model performance, focusing on those designed to address class imbalance: Standard Binary Cross-Entropy Loss (SLF), Focal Loss with Dynamic Scaling (FLF), Weighted Binary Cross-Entropy Loss (WLF), and Weighted Binary Cross-Entropy Loss with Reduced Negative Samples (RLF). Experiments were conducted in imbalanced datasets and details of these loss functions could be found in Appendix Section E. As shown in Table 3, the RLF loss function consistently outperformed the others across four benchmark imbalanced datasets, demonstrating its effectiveness for imbalanced DTI prediction. Although RLF did not always yield the highest scores on individual metrics, it showed stable and robust performance. By incorporating a weighting parameter $\varpi$ to reduce the weight of negative samples while maintaining the positive sample weight, RLF addressed class imbalance while preserving the model’s discriminative ability. These results, observed across four distinct imbalanced datasets, further validated RLF’s strong generalization capability.